OpenStudy (anonymous):

factor completely m^6-216n^6 i want to know if i did it right..thanks

OpenStudy (anonymous):

i get (m^2-6n^2)(m^4+6n^2m^2+36n^4)

OpenStudy (jhannybean):

you are correct

OpenStudy (markchernioglo):

That looks correct to me lordofpens. :)

OpenStudy (jhannybean):

This is a difference of cubes,

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

how about 121a^2-(3b^3-5x)^2

OpenStudy (jhannybean):

\[m^6 -216n^6 = (m^2)^3-(6n^2)^3\]\[a^3-b^3 = (a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

i get (11a-3b63+5x)(11a+3b^3-5x)

OpenStudy (jhannybean):

Idk what you're doing there, but your first answer was correct :)